# Operator Growth in Disordered Spin Chains: Indications for the Absence of Many-Body Localization

### v3

@@ IV.1 Non-interacting case
$\partial_{t}\sigma^{z}_{i}=-i[\sigma^{z}_{i},H]=-(j_{i}-j_{i-1})\,.$ (IV.5)
We note, furthermore, that $\overline{|h_{j+1}-h_{j}|}=2/3$, i.e., the factor $\bar{J}^{(0)}_{j,j+1}/|h_{j}-h_{j+1}|$ is small on average if the disorder strength $D$ is large. However, if two neighboring sites have roughly the same potential, then this factor can become large or might even diverge. Rare resonances thus might destroy the perturbative construction if they proliferate. An example how such resonances can affect the transformed Hamiltonian in shown in Fig. 15, bottom left panel. We will come back to this point below.
@@
$T^{(k)}=\frac{1}{2}\sum_{j}\frac{\bar{J}^{(k)}_{j,j+k+1}}{h^{(k)}_{j+k+1}-h^{(k)}_{j}}(\sigma^{+}_{j}\sigma^{z}_{j+1}\dots\sigma^{z}_{j+k}\sigma^{-}_{j+k+1}-h.c.)$ (IV.8)
with $k=1$. This will again, to leading order, renormalize the local fields and generate a third-nearest neighbor hopping which again can be eliminated to leading order by a the Schrieffer-Wolff transformation (IV.8) with $k=2$. I.e., we can define a series of successive Schrieffer-Wolff transformations which lead to longer and longer range interactions which are, however, suppressed as $D^{-2k}$. At the same time, they define an exponentially converging series for the renormalized random fields $h^{(k)}_{j}$. If there are no resonances, then we can drop the hopping term when all the amplitudes $J^{(k)}$ have fallen below some threshold, giving us an effective Hamiltonian with renormalized magnetic fields. In this approximation, which can be justified at very strong disorder and which corresponds to summing up exactly certain classes of diagrams, the conserved local charges are just the $\sigma^{z}_{j}$ themselves but with renormalized amplitudes. This can be seen as a consistency check that at very large disorder such an effective description exists because the hopping is an irrelevant perturbation.
@@
IV.2 Interacting case
The interacting case, on the other hand, is completely different and none of the above arguments hold. We are, however, able to generalize Eq. (IV.4) and to obtain the exact Schrieffer-Wolff transformation for the XXZ chain and arbitrary interaction strength $\Delta$. We find
$T$ $=$ $\sum_{j}\frac{\bar{J}^{(0)}_{j,j+1}(\sigma^{+}_{j}\sigma^{-}_{j+1}-h.c.)}{2(h_{j}-h_{j+1})(2\bar{\Delta}^{2}-2(h_{j}-h_{j+1})^{2})}$ (IV.9)
@@ IV.2 Interacting case
If we stop the transformation (IV.3) at first order, the transformed Hamiltonian $H^{(1)}$ already has $29N$ distinct operators if the chain length $N$ is sufficiently large. This includes terms acting on $5$ neighboring sites but also terms such as $\sigma^{z}_{j}\sigma^{z}_{j+1}$ and $\sigma^{z}_{j}\sigma^{z}_{j+2}\sigma^{z}_{j+3}$. Even for very strong disorder, the transformation thus does not just lead to a renormalization of the random fields but necessarily leads to a broadening of the diagonal basis. Symbolically, we can also construct $T^{(1)}$ to eliminate the off-diagonal terms in $H^{(1)}$. We find that $T^{(1)}$ already consists of $786N$ distinct operators. The transformed Hamiltonian $H^{(2)}$ then has $30618N$ many terms. I.e., even if we stop the Schrieffer-Wolff transformation at lowest order, the number of terms in $T^{(k)}$ and $H^{(k)}$ is growing exponentially. Furthermore, there is no distinct set of diagrams which we can simply sum up to all orders. Exponentially many new diagrams are created at each order.
@@
The related and most fundamental problem, however, is that lattice sites $j$ and $j+\ell$ a distance $\ell$ apart will be coupled by $\sim 3^{2}4^{\ell-1}$ distinct operators because on each site we can put in principle one of the operators $\{\sigma^{x},\sigma^{y},\sigma^{z},\mathbb{I}\}$ but the string of operators of length $\ell+1$ has to start and end with one of the three Pauli matrices. While not all of these combinations of operators will be realized in a microscopic model such as the XXX chain, there will be exponentially many distinct operators which couple these two sites, see Fig. 15. Thus there is, in general, no reason to expect this perturbative construction to work because the overall effective amplitude of terms coupling these two sites will be $\bar{J}_{j,j+\ell}\sim D^{-\ell}4^{\ell}\sim\mathcal{O}(1)$. I.e., there is no small parameter anymore. This reasoning is supported by the results of symbolic calculations shown in Fig. 15. Here we have calculated the sum $\tilde{s}_{\ell}^{(1)}(k)$ of the amplitudes of all terms obtained in the Schrieffer-Wolff transformation (IV.9) to order $k$ which have support on $\ell$ sites, i.e., which connect sites a distance $\ell-1$ apart. For a periodic chain with even length $N$, there are $\sim N$ distinct pairs of sites which are a distance $21,\cdots,N/$ 2 apart. I.e., for a specific pair the overall coupling is on average $\sim\tilde{s}_{\ell}^{(1)}(k)/N$ and, as can be seen in Fig. 15, is in general not exponentially decreasing with $\ell$. That the number of terms generated in such a transformation with support on $\ell$ sites is exponentially increasing with $\ell$ makes it very questionable that a construction of local conserved charges in this perturbative manner is a controlled approximation. This is a major issue even before thinking about the problem of resonances and thermal inclusions (regions for a given disorder configuration which, by chance, are only weakly disordered). It seems to us that the exponential increase of the number of terms in such transformations is a point which has not been properly addressed in the proof of MBL in Refs. [47, 48] where the author seems to be only concerned about the effect of resonances. While it is not impossible that for disorder configurations without resonances and thermal inclusions the perturbative construction of local charges is working in the thermodynamic limit—for example, because of interference effects between different terms coupling the same sites—any proof needs to provide rigorous arguments that the exponentially many terms do not overwhelm the exponential small parameter in the expansion. Such arguments seem to be missing so far. We note that for a finite system the construction will always eventually converge because the finite length provides a cutoff and $\tilde{s}_{\ell}^{(1)}(k)$ will eventually drop off exponentially with $\ell$.
@@
The destabilizing effects of resonances and thermal inclusions have received considerable attention in the literature already [56, 57, 42]. From the perspective of perturbative Schrieffer-Wolff transformations, they lead to the following additional problems: The renormalized couplings for an operator acting on $\ell$ lattice sites will in its denominator typically contain a linear combination of all the random fields on these $\ell$ sites. Given that there are exponentially many such operators, the simple scaling argument for the non-proliferation of resonances in the Anderson case is clearly no longer applicable in the interacting case. This can be seen in Fig. 15, bottom left panel, where the sum of the amplitudes of terms in the Hamiltonian with support on one site, $\tilde{s}_{1}(k)$, in the interacting case is large even for samples which do not have nearest-neighbor resonances. Thermal inclusions—regions of the lattice which have similar potentials—lead, furthermore to an additional destabilization of the perturbative transformation. This can be seen, for example, in symbolic calculations of the unitary transformation (IV.1) with the operator (IV.9) to higher orders. In configurations where such regions exist, the transformation generates long-range terms with large amplitudes even if the disorder $D$ is extremely strong. If, on the other hand, we exclude configurations with such regions and with any pairwise resonances by hand, then the amplitudes of any individual operator typically decay with the length of its support. Another manifestation of this issue can be seen when performing $k$ successive lowest order Schrieffer-Wolff transformations for a finite chain of length $N$. We want to stop the transformations again once the 1-norm of the off-diagonal part is below some threshold $\varepsilon$. In contrast to the non-interacting case, we now have exponentially many terms connecting sites a given distance apart. If we assume that individual amplitudes in the off-diagonal part are suppressed as $D^{-2k}$, then the condition for stopping the transformations becomes $\text{e}^{N}D^{-2k}<\varepsilon$ and we would thus expect that we need $k\sim N$ steps for $N$ sufficiently large. This is, however, not what we observe for typical disorder configurations. The results of our symbolic calculations point instead to an increase of the number of transformations $k\sim\text{e}^{N}$. This implies that some amplitudes in the off-diagonal part are $\sim\mathcal{O}(1)$ and that we have to eliminate almost all of the off-diagonal terms to achieve convergence.
@@
In summary, the iterative transformation of the Hamiltonian cannot simply be transferred from the non-interacting to the interacting case. There are two main issues: Even if one considers only a lowest order transformation, one is not dealing with a constant number of operators but rather with an exponentially increasing number of operators, both in the Schrieffer-Wolff transformation $T^{(k)}$ as well as in the transformed Hamiltonian $H^{(k)}$. Even if one tries to only perform a single unitary transformation (IV.1) with the operator $T$, the number of distinct operators in the transformed Hamiltonian is increasing exponentially with the support of the operator which thus can overwhelm the exponential decay of the amplitude of individual operators with the length of their support. We note that different but related arguments for the failure of a perturbative construction of conserved charges in the interacting case have recently been made in Ref. [58]. Furthermore, regions with similar lattice potentials result in individual operators with large support which, nevertheless, have large amplitudes even for very strong disorder. The latter issue is related to the well-known avalanche instability [56, 57, 59, 42].
V Conclusions
@@
We have considered operator growth in one-dimensional lattice models and its connection to localization from two different perspectives. First, we studied the multiple commutator $[H,A]^{(k)}$, which appears in the Euclidean time evolution, for a local operator $A$ and a Hamiltonian $H$ which is a sum of nearest-neighbor terms. For a generic, ergodic system, we have shown that the number of distinct terms generated is exponentially increasing with the order $k$ of the commutator. This leads to a strict bound for the commutator norm $s(k)$ which grows faster than exponential with $k$, a result which was first demonstrated in Ref. [30]. Since this result is based on a simple counting of connected clusters generated by the commutator, it is expected that this bound is asymptotically tight for an ergodic system. This operator growth hypothesis is challenged by results for systems such as the random transverse Ising chain which are supposed to have a localized phase but, nevertheless, always show almost factorial operator growth. Here we systematically addressed this issue. We started by providing strict bounds for two special cases. First, we considered the Hamiltonian $H$ of a free-fermion system. In this case, all the terms generated by the commutator of the Hamiltonian with a one-body operator $A$ will also always be one-body operators. For a short-range Hamiltonian, the spatial range of operators generated at order $k$ will be proportional to $k$. The number of distinct terms in the commutator can therefore only grow $\sim k^{2}$ as compared to the exponential growth in the general, interacting case. We also proved an exponential bound for the commutator norm in this case. Second, we considered a localized system, either interacting or non-interacting, and showed that an exponential bound for $s(k)$ exists in this case as well. Physically, this can be understood as follows: a localized system has a set of local conserved charges and the Hamiltonian can be written in terms of these charges. Commuting the Hamiltonian with a local operator $A$ then only involves those charges localized near the site the operator $A$ is acting on. Thus, up to exponential tails, only operators with support on a number of sites which is less than the localization length will significantly contribute to the norm. We note that an almost factorial versus an exponential growth of the commutator norm also leads to distinctly different Lieb-Robinson bounds and asymptotic behavior of matrix elements in the energy eigenbasis. This is further discussed in App. B and App. C.
To distinguish a localized interacting or non-interacting system from a non-localized free fermion system, it is important to look beyond the overall norm which has an exponential bound in both cases. Here we have shown that for a non-localized free-fermion system at large commutator order $k$, the norm $s_{l}(k)$ of terms which have support on $l$ sites will grow exponentially with $k$ and is independent of the length of the support $l$. In contrast, the norm scales as $s_{l}(k)\sim\exp(k)\exp(-l)$ for a localized system. I.e., at a fixed order $k$, the contribution of terms in the commutator is exponentially suppressed with the length of their support $l$.
We thus have proven sharp criteria to distinguish between free-fermionic localized and non-localized and interacting localized and non-localized systems based on the growth of the total commutator norm $s(k)$ and the contributions $s_{l}(k)$ to this norm by operators with support on $l$ sites. To illustrate our results, we then considered specific microscopic examples, namely the XX chain (nearest-neighbor free-fermion model), the Anderson model, the Aubry-André model, and the XXX (Heisenberg) chain with and without quenched disorder. For the XX case we have demonstrated that a complete, exact result for the commutator $[H,\sigma^{z}_{0}]^{(k)}$ to all orders $k$ can be derived. These exact results confirmed and illustrated our general results derived earlier. For the Anderson model, we derived an exponential bound for the total norm which does depend on the disorder strength $D$. Based on a graph theoretical approach, we also showed that $s_{l}(k)\sim\exp(-l)$ for fixed $k$. I.e., the contribution of terms in the commutator to the total norm is indeed exponentially suppressed with the length $l$ of their support. As another example, we considered the Aubry-André model which has a transition from a non-localized to a localized phase at some finite strength of the applied incommensurate lattice potential. Importantly, we could demonstrate that the non-localized and localized phases of the model can clearly be distinguished by considering the norm $s_{l}(k)$.
Having checked our results for non-interacting cases, we then turned to the Heisenberg (XXX) chain with quenched disorder. This model has been suggested to have a transition from an ergodic to a many-body localized phase at some finite disorder strength $D$. In particular, for the MBL phase an effective Hamiltonian written in terms of localized conserved charges is supposed to exist as in the Anderson case but with couplings between these charges which are exponentially decaying with distance. If true, then our proof for localized phases applies. Thus $s(k)$ can grow at most exponentially in such a phase and $s_{l}(k)\sim\exp(-l)$ at fixed $k$. We used symbolic calculations of the commutator and were able to reach order $k=16$ at which point we are considering terms located on a lattice of $332k+1=$ 33 sites. This is substantially larger than what can be done using exact diagonalizations. We found that the growth of the norm of the commutator $s(k)$ appears to be faster than exponential and consistent with the generic, ergodic bound for non-localized systems for all considered disorder strengths up to $D=25$. Similarly, we find no indication that $s_{l}(k)\sim\exp(-l)$. To the contrary, we find that $s_{l}(k)/s_{1}(k)$ has a maximum which appears to shift to larger and larger $l$ with increasing $k$. At the very least, we claim that our results clearly demonstrate that there is no convincing case for an MBL phase for system sizes up to $33$ sites and disorder strengths up to $D=25$.
@@
As a second approach, we considered unitary Schrieffer-Wolff transformations of the entire Hamiltonian to perturbatively eliminate the hopping terms and construct local conserved charges. We showed that such an approach will succeed in the Anderson case because the Hamiltonian in this case consists of one-body operators and this property is not changed by unitary transformations. Consequently, there are only two terms possible in the effective Hamiltonian which can couple two sites $j$ and $k$: $c^{\dagger}_{j}c_{k}$ and $c_{j}c_{k}^{\dagger}$. This allows to define consecutive lowest order Schrieffer-Wolff transformations and to sum up the entire series. In this approximation, valid at very strong disorder, the local conserved charges are simply the $\sigma^{z}_{j}$ operators with renormalized fields. The longer-range hoppings generated by the transformations have amplitudes which decay exponentially with the range of the hopping process. The minimum gap between random fields on a cluster of length $\ell$ on the other hand—which enters the effective hopping in the denominator—scales as $\ell^{-2}$ and thus cannot overcome the general suppression of the hopping process which scales as $D^{-2\ell}$. This means, in particular, that resonances cannot proliferate.
@@
All of this changes drastically once interactions are included. In particular, there are then $\sim 4^{\ell}$ different terms which can connect sites a distance $\ell$ apart. We confirm this exponential growth with the length of the support by symbolic calculations. A priori, there is thus no reason to believe that a perturbative construction of the local conserved charges by unitary transformations is still possible in the interacting case. Even if we assume that there are no resonances or thermal inclusions (regions with random fields of similar strengths), there is a fundamental problem with this construction: while the amplitude of an individual term connecting sites a distance $\ell$ apart will still be $\sim D^{-2\ell}$, there are exponentially many such terms so that the total effective amplitude is $\sim 4^{\ell}D^{-2\ell}\sim\mathcal{O}(1)$. The small parameter in the perturbative construction is lost. Our symbolic calculations do confirm this issue and indeed show that the total effective amplitude is not decaying with the support of the operator in contrast to the non-interacting case. As far as we can tell, this issue is not addressed at all in the proof of MBL by Imbrie [47, 48] which uses the same kind of unitary transformations but seems to be only concerned about resonances. The latter is also a much more severe problem in the interacting case. Even if we assume that the perturbative construction is convergent for carefully chosen disorder configurations without resonances and thermal inclusions due to some miraculous cancellation of contributions stemming from different commutator paths, many-body resonances will destroy the approach when present. A long-range term typically involves operators on all sites between the left most and the right most site. Consequently, the renormalization will involve linear combinations of all the random fields on this cluster and there are exponentially many of such terms. Clearly, the argument for the non-proliferation of resonances made in the non-interacting case is then no longer valid. At the very least, we believe that we have shown that there is no reason to believe that the proof in Refs. [47, 48] is valid. It seems to ignore that, in general, the number of terms connecting distant sites grows exponential with the distance between the sites. No argument is made why the construction remains controlled by a small parameter despite the exponential number of terms and why resonances do not proliferate. Our symbolic calculations for finite chains and finite orders of the transformation suggest that such an argument, in fact, cannot be made and that a perturbative construction of local conserved charges in this way is impossible except for $D=\infty$ where the model is trivially localized.
While we have concentrated here on the Heisenberg chain with quenched disorder, we note that in the attempted proof of MBL by Imbrie [47, 48] the transverse Ising chain with random couplings and fields in considered instead. For the latter model, it has recently been rigorously proven [37] that the norm of a local operator when commuted $k$-times with the Hamiltonian grows faster than exponential with $k$ and, in fact, asymptotically like the general bound (II.12). Here, we have shown that if a local operator remains quasi-local when commuted with a Hamiltonian then the total norm cannot grow faster than exponential. If we negate this statement, then this means that if the commutator norm of a local operator with a Hamiltonian grows faster than exponential — as in the transverse Ising case — then the commutator is necessarily not a quasi-local operator. It can at best be algebraically localized. I.e., our results come close to proving that the operator growth hypothesis is valid in the strict sense that an almost factorial growth indicates an ergodic systems and that an MBL phase does not exist in the transverse Ising chain. The only assumption we have made is that for the commutator to remain quasi-local, contributions to the connected clusters coming from sites far away from the local operator are exponentially suppressed, see App. A. This seems to be a physical requirement for a localized model. Mathematically it is also hard to see how a local operator can otherwise remain quasi-local. This would then require some miraculous cancellation between different paths which has to happen for any local operator and most disorder configurations. For the transverse Ising chain we already know for sure that such destructive interference effects between different paths do not exist because a gauge can be chosen such that all the terms in the commutator are positive [37]. Making these arguments even more rigorous might be the last remaining step in proving that localization is not possible in general in the many-body case.

### v4

@@ VI.1 Non-interacting case
$\partial_{t}\sigma^{z}_{i}=-i[\sigma^{z}_{i},H]=-(j_{i}-j_{i-1})\,.$ (VI.5)
We note, furthermore, that $\overline{|h_{j+1}-h_{j}|}=2/3$, i.e., the factor $\bar{J}^{(0)}_{j,j+1}/|h_{j}-h_{j+1}|$ is small on average if the disorder strength $D$ is large. However, if two neighboring sites have roughly the same potential, then this factor can become large or might even diverge. Rare resonances thus might destroy the perturbative construction if they proliferate. An example how such resonances can affect the transformed Hamiltonian is shown in Fig. 18, bottom left panel. We will come back to this point below.
@@
$T^{(k)}=\frac{1}{2}\sum_{j}\frac{\bar{J}^{(k)}_{j,j+k+1}}{h^{(k)}_{j+k+1}-h^{(k)}_{j}}(\sigma^{+}_{j}\sigma^{z}_{j+1}\dots\sigma^{z}_{j+k}\sigma^{-}_{j+k+1}-h.c.)$ (VI.8)
with $k=1$. This will again, to leading order, renormalize the local fields and generate a third-nearest neighbor hopping which again can be eliminated to leading order by a the Schrieffer-Wolff transformation (VI.8) with $k=2$. I.e., we can define a series of successive Schrieffer-Wolff transformations which lead to longer and longer range interactions which are, however, suppressed as $D^{-2k}$. At the same time, they define an exponentially converging series for the renormalized random fields $h^{(k)}_{j}$. If there are no resonances, then we can drop the hopping term when all the amplitudes $J^{(k)}$ have fallen below some threshold, giving us an effective Hamiltonian with renormalized magnetic fields. In this approximation, which can be justified at very strong disorder and which corresponds to summing up exactly certain classes of diagrams, the conserved local charges are just the $\sigma^{z}_{j}$ themselves but with renormalized amplitudes. This can be seen as a consistency check that at very large disorder such an effective description exists because the hopping is an irrelevant perturbation.
@@
VI.2 Interacting case
The interacting case, on the other hand, is completely different and none of the above arguments hold. We are, however, able to generalize Eq. (VI.4) and to obtain the exact Schrieffer-Wolff transformation for the XXZ chain and arbitrary interaction strength $\Delta$. We find
$T$ $=$ $\sum_{j}\frac{\bar{J}^{(0)}_{j,j+1}(\sigma^{+}_{j}\sigma^{-}_{j+1}-h.c.)}{2(h_{j}-h_{j+1})(2\bar{\Delta}^{2}-2(h_{j}-h_{j+1})^{2})}$ (VI.9)
@@ VI.2 Interacting case
If we stop the transformation (VI.3) at first order, the transformed Hamiltonian $H^{(1)}$ already has $29N$ distinct operators if the chain length $N$ is sufficiently large. This includes terms acting on $5$ neighboring sites but also terms such as $\sigma^{z}_{j}\sigma^{z}_{j+1}$ and $\sigma^{z}_{j}\sigma^{z}_{j+2}\sigma^{z}_{j+3}$. Even for very strong disorder, the transformation thus does not just lead to a renormalization of the random fields but necessarily leads to a broadening of the diagonal basis. Symbolically, we can also construct $T^{(1)}$ to eliminate the off-diagonal terms in $H^{(1)}$. We find that $T^{(1)}$ already consists of $786N$ distinct operators. The transformed Hamiltonian $H^{(2)}$ then has $30618N$ many terms. I.e., even if we stop the Schrieffer-Wolff transformation at lowest order, the number of terms in $T^{(k)}$ and $H^{(k)}$ is growing exponentially. Furthermore, there is no distinct set of diagrams which we can simply sum up to all orders. Exponentially many new diagrams are created at each order.
@@
The related and most fundamental problem, however, is that lattice sites $j$ and $j+\ell$ a distance $\ell$ apart will be coupled by $\sim 3^{2}4^{\ell-1}$ distinct operators because on each site we can put in principle one of the operators $\{\sigma^{x},\sigma^{y},\sigma^{z},\mathbb{I}\}$ but the string of operators of length $\ell+1$ has to start and end with one of the three Pauli matrices. While not all of these combinations of operators will be realized in a microscopic model such as the XXX chain, there will be exponentially many distinct operators which couple these two sites, see Fig. 18. Thus there is, in general, no reason to expect this perturbative construction to work because the overall effective amplitude of terms coupling these two sites will be $\bar{J}_{j,j+\ell}\sim D^{-\ell}4^{\ell}\sim\mathcal{O}(1)$. I.e., there is not necessarily a small parameter anymore. This reasoning is supported by the results of symbolic calculations shown in Fig. 18. Here we have calculated the sum $\tilde{s}_{\ell}^{(1)}(k)$ of the amplitudes of all terms obtained in the Schrieffer-Wolff transformation (VI.9) to order $k$ which have support on $\ell$ sites, i.e., which connect sites a distance $\ell-1$ apart. For a periodic chain with even length $N$, there are $\sim N$ distinct pairs of sites which are a distance $21,\cdots,N/$ 2 apart. I.e., for a specific pair the overall coupling is on average $\sim\tilde{s}_{\ell}^{(1)}(k)/N$ and, as can be seen in Fig. 18, is in general not exponentially decreasing with $\ell$. That the number of terms generated in such a transformation with support on $\ell$ sites is exponentially increasing with $\ell$ makes it very questionable that a construction of local conserved charges in this perturbative manner is a controlled approximation. This is a major issue even before addressing the issue of resonances and thermal inclusions (regions for a given disorder configuration which, by chance, are only weakly disordered). It seems to us that the exponential increase of the number of terms in such transformations is a point which has not been properly addressed in the proof of MBL in Refs. [48, 49] where the author seems to be only concerned about the effect of resonances. While it is not impossible that for disorder configurations without resonances and thermal inclusions the perturbative construction of local charges is working in the thermodynamic limit for very strong disorder, any proof needs to provide rigorous arguments that the exponentially many terms do not overwhelm the exponential small parameter in the expansion. Such arguments seem to be missing so far. We note that for a finite system the construction will always eventually converge because the finite length provides a cutoff and $\tilde{s}_{\ell}^{(1)}(k)$ will eventually drop off exponentially with $\ell$.
@@
The destabilizing effects of resonances and thermal inclusions have received considerable attention in the literature already [57, 58, 43]. From the perspective of perturbative Schrieffer-Wolff transformations, they lead to the following additional problems: The renormalized couplings for an operator acting on $\ell$ lattice sites will in its denominator typically contain a linear combination of all the random fields on these $\ell$ sites. Given that there are exponentially many such operators, the simple scaling argument for the non-proliferation of resonances in the Anderson case is clearly no longer applicable in the interacting case. This is confirmed by the symbolic calculations shown in Fig. 18, bottom left panel, where the sum of the amplitudes of terms in the Hamiltonian with support on one site, $\tilde{s}_{1}(k)$, in the interacting case is large even for samples which do not have nearest-neighbor resonances. Thermal inclusions—regions of the lattice which have similar potentials—lead, furthermore to an additional destabilization of the perturbative transformation. This can be seen, for example, in symbolic calculations of the unitary transformation (VI.1) with the operator (VI.9) to higher orders. In configurations where such regions exist, the transformation generates long-range terms with large amplitudes even if the disorder $D$ is extremely strong. If, on the other hand, we exclude configurations with such regions and with any pairwise resonances by hand, then the amplitudes of any individual operator typically decay with the length of its support. Another manifestation of this issue can be seen when performing $k$ successive lowest order Schrieffer-Wolff transformations for a finite chain of length $N$. We want to stop the transformations again once the 1-norm of the off-diagonal part is below some threshold $\varepsilon$. In contrast to the non-interacting case, we now have exponentially many terms connecting sites a given distance apart. If we assume that individual amplitudes in the off-diagonal part are suppressed as $D^{-2k}$, then the condition for stopping the transformations becomes $\text{e}^{N}D^{-2k}<\varepsilon$ and we would thus expect that we need $k\sim N$ steps for $N$ sufficiently large. This is, however, not what we observe for typical disorder configurations. The results of our symbolic calculations point instead to an increase of the number of transformations $k\sim\text{e}^{N}$. This implies that some amplitudes in the off-diagonal part are $\sim\mathcal{O}(1)$ and that we have to eliminate almost all of the off-diagonal terms to achieve convergence.
@@
In summary, the iterative transformation of the Hamiltonian cannot simply be transferred from the non-interacting to the interacting case. There are two main issues: Even if one considers only a lowest order transformation, one is not dealing with a constant number of operators but rather with an exponentially increasing number of operators, both in the Schrieffer-Wolff transformation $T^{(k)}$ as well as in the transformed Hamiltonian $H^{(k)}$. Even if one tries to only perform a single unitary transformation (VI.1) with the operator $T$, the number of distinct operators in the transformed Hamiltonian is increasing exponentially with the support of the operator which thus can overwhelm the exponential decay of the amplitude of individual operators with the length of their support. We note that different but related arguments for the failure of a perturbative construction of conserved charges in the interacting case have recently been made in Ref. [59]. Furthermore, regions with similar lattice potentials result in individual operators with large support which, nevertheless, have large amplitudes even for very strong disorder. The latter issue is related to the well-known avalanche instability [57, 58, 60, 43].
VII Conclusions
@@
We have considered operator growth in one-dimensional lattice models and its connection to localization from two different perspectives. First, we studied the multiple commutator $[H,A]^{(k)}$, which appears both in real-time and Euclidean time evolution, for a local operator $A$ and a Hamiltonian $H$ which is a sum of nearest-neighbor terms. For a generic, ergodic system, we have shown that the number of distinct terms generated is exponentially increasing with the order $k$ of the commutator. This leads to a strict bound for the commutator norm $s(k)$ in nearest-neighbor models which grows almost factorially with $k$, a result which was first demonstrated in Ref. [30]. Since this result is based on a simple counting of connected clusters generated by the commutator, it is expected that this bound is asymptotically tight for an ergodic system. This operator growth hypothesis is challenged by recent results for systems such as the random field Ising chain which are supposed to have a localized phase but, nevertheless, always show almost factorial operator norm growth independent of the strength of the disorder. Here we have systematically addressed this issue. We started by analytically obtaining strict upper bounds for the operator norm for free fermion and for localized systems. For a free fermion system, all the terms generated by the commutator of the Hamiltonian $H$ with a one-body operator $A$ will also always be one-body operators. For a short-range Hamiltonian, the spatial range of operators generated at order $k$ will be proportional to $k$. The number of distinct terms in the commutator can therefore only grow $\sim k^{2}$ as compared to the exponential growth in the general, interacting case. This leads to a rigorous exponential upper bound for the commutator norm in this case. Second, we considered the case of strict exponential localization where we demanded that a local operator $A$ remains exponentially localized when commuted with $H$ up to exponential tails. We proved that in this case a strict upper exponential norm bound does exist as well. The almost factorial norm growth proven for the Ising model with random fields for all disorder strengths and strongly suggested for the random field Heisenberg model by symbolic calculations in this work is thus inconsistent with a strict exponential localization of operators. We then considered the alternative definition of localization based on the unitary transformation $U$ of a microscopic Hamiltonian $H$ to an effective Hamiltonian $\tilde{H}$ which has an infinite set of local conserved charges. Importantly, local operators in $H$ are mapped by $U$ onto quasi-local operators. For the non-interacting case we proved that this definition is equivalent to the strict exponential localization of local operators. The contribution of terms in the commutator with support on $l$ sites to the norm does scale as $s_{l}(k)\sim\exp(k)\exp(-l)$. For the interacting case, however, we were only able to prove a weaker bound $s_{l}(k)\sim l^{k}\exp(-l)$ which is consistent with an almost factorial growth of the total norm. Furthermore, this bound means that for very large $k$, $s_{l}(k)$ grows in exactly the same way as in a generic ergodic system. The question then becomes whether the transformation $U$ exists and models such as the random field Ising and Heisenberg models have a non-ergodic phase in which, however, local operators eventually completely delocalize or if the transformation $U$ simply does not exist. We have also shown that even if the quasi-local unitary $U$ does exist, the spreading of local operators through the entire lattice is generically expected to lead to transport of globally conserved charges so that such phases, if they do exist, should not be called many-body localized.
@@
To further study the existence of the transformation $U$, we considered unitary Schrieffer-Wolff transformations of the XXZ chain to perturbatively eliminate the hopping terms and construct $U$. We showed that such an approach will succeed in the Anderson case because the Hamiltonian in this case consists of one-body operators and this property is not changed by unitary transformations. Consequently, there are only two terms possible in the effective Hamiltonian which can couple two sites $j$ and $k$: $c^{\dagger}_{j}c_{k}$ and $c_{j}c_{k}^{\dagger}$. This allows to define consecutive lowest order Schrieffer-Wolff transformations and to sum up the entire series. In this approximation, valid at very strong disorder, the local conserved charges are simply the $\sigma^{z}_{j}$ operators with renormalized fields. The longer-range hoppings generated by the transformations have amplitudes which decay exponentially with the range of the hopping process. The minimum gap between random fields on a cluster of length $\ell$ on the other hand—which enters the effective hopping in the denominator—scales as $\ell^{-2}$ and thus cannot overcome the general suppression of the hopping process which scales as $D^{-2\ell}$. This means, in particular, that resonances cannot proliferate.
@@
All of this changes drastically once interactions are included. In particular, there are then $\sim 4^{\ell}$ different terms which can connect sites a distance $\ell$ apart. We confirm this exponential growth with the length of the support by symbolic calculations. A priori, there is thus no reason to believe that a perturbative construction of the local conserved charges by unitary transformations is still possible in the interacting case. Even if we assume that there are no resonances or thermal inclusions (regions with random fields of similar strengths), there is potentially a fundamental problem with this construction: while the amplitude of an individual term connecting sites a distance $\ell$ apart will still be $\sim D^{-2\ell}$, there are exponentially many such terms so that the total effective amplitude is $\sim 4^{\ell}D^{-2\ell}\sim\mathcal{O}(1)$. The small parameter in the perturbative construction is in general lost even before considering the issue of resonances. While this simple scaling argument does not preclude that the perturbative construction in the absence of resonances can succeed for very strong disorder such that $4^{\ell}D^{-2\ell}\ll 1$ our symbolic calculations show that the total effective amplitude is not decaying with the support of the operator in contrast to the non-interacting case. As far as we can tell, this issue is not addressed in the proof of MBL by Imbrie [48, 49] which uses the same kind of unitary transformations but seems to be only concerned about resonances. The latter is also a much more severe problem in the interacting case. Even if we assume that the perturbative construction is convergent without resonances and thermal inclusions, many-body resonances will destroy the approach when present. A long-range term typically involves operators on all sites between the left most and the right most site. Consequently, the renormalization will involve linear combinations of all the random fields on this cluster and there are exponentially many of such terms. Clearly, the argument for the non-proliferation of resonances made in the non-interacting case is then no longer valid. At the very least, we believe that we have shown that there are certain aspects of the perturbative proof in Refs. [48, 49] which would need to be addressed to make it more convincing. In particular, the discussion of the construction seems to ignore that, in general, the number of terms connecting distant sites grows exponential with the distance between the sites. It is not obvious why the construction remains controlled by a small parameter despite the exponential number of terms and why resonances do not proliferate. Our symbolic calculations for finite chains and finite orders of the transformation suggest that such an argument, in fact, cannot be made and that a perturbative construction of local conserved charges in this way is impossible except for $D=\infty$ where the model is trivially localized. We therefore conclude that we have found strong evidence that the unitary $U$ does not exist and that the almost factorial growth of the operator norm for the random field Ising and Heisenberg models indicates that these models are ergodic in the thermodynamic limit for all disorder strengths, consistent with the operator growth hypothesis. In a generic scenario, this implies that these systems in the thermodynamic limit and at very long times will ultimately exhibit finite transport coefficients.
As an outlook, we note that under the assumption that the unitary transformation $U$ exists, we have only been able to prove the bound $s_{l}(k)\sim l^{k}\exp(-l)$ which is consistent with an almost factorial growth of the total norm, $s(k)\sim(k/\ln k)^{k}$. However, we have not used any specific properties of a microscopic model such as the random Ising or Heisenberg model so it might be possible that a stricter bound for $s_{l}(k)$ can be derived for special cases which leads to a parametrically slower growth than $s(k)\sim(k/\ln k)^{k}$ and is thus in direct contradiction with the rigorous result for the norm growth from Ref. [37]. This seems to be a promising direction for future research.